An English reading list has 10 American novels and 7 English novels. A student must read 4 from the list, and at least 2 must be American novels. In how many different ways can the 4 books be selected?
Answer by greenestamps(13203)



10 American and 7 English; choose 4, with at least 2 of them American.

2 American and 2 English: (10 choose 2) times (7 choose 2) = 45*21 = 945

3 American and 1 English: (10 choose 3) times (7 choose 1) = 120*7 = 840

4 American and 0 English: (10 choose 4) times (7 choose 0) = 210*1 = 210

Total: 945+840+210 = 1995

ANSWER: 1995